OpenStudy (owlcoffee):

Need some help with the analytic study and graphic representation of the following function:

OpenStudy (owlcoffee):

\[f:f(x)=(x-2)e ^{3x-1}\]

OpenStudy (anonymous):

What exactly do you need?

OpenStudy (owlcoffee):

First I want to know the limits of that funtion as x approaches + and - inifinity. I got that both gave me infinity, but I believe I did something wrong there:

OpenStudy (anonymous):

How did you find your limits?

OpenStudy (owlcoffee):

\[\lim_{x \rightarrow +\infty}(x-2)e ^{3x-1}=+\infty \]

OpenStudy (anonymous):

what made you conclude that?

OpenStudy (anonymous):

Happens to be that you're right. Just want to hear your argument.

OpenStudy (owlcoffee):

Well, I know that any number multiplied by infinity is equal to inifnity, and if I substract -1 it'll still be infinity, so I would end un havinf e^inf wich is equal to inifnity. (x-2), if I replace infinity there, I would end up having infinity, so the total would be: (infinity)(infinity) = infinity, and since I'm doing it from the right... +infinity,

OpenStudy (anonymous):

agree. What about as x------> - infinity?

OpenStudy (owlcoffee):

That's where my problem starts: So if I replace infinity it would be 0?

OpenStudy (anonymous):

It does approach 0.

OpenStudy (anonymous):

as x approaches negative infinity, e^(3x-1) approaches 0.

OpenStudy (anonymous):

An exponential grows facster than a polynomial (x-2).

OpenStudy (anonymous):

faster*

OpenStudy (anonymous):

so that entire expression approaches 0.

OpenStudy (owlcoffee):

So that's because the exponent gets negative, and because of the property: \[a ^{-n}=\frac{ 1 }{a ^{n} }\] and because the e grows faster to infinity than the polynomial, it's 0?

OpenStudy (anonymous):

yes. Look at e^x...as x------> negative infinity, e^x approaches 0...an it will outrun any polynomial as x approahes negative infinity.

OpenStudy (anonymous):

e^x and e^(3x-1) is the same function when you want to look at end-behavior as x approaches negative infinity.

OpenStudy (owlcoffee):

nice... Now I would need to study the growth of that function, meaning that I have to derivate it and study the behavior of the slopes.

OpenStudy (anonymous):

I'm not clear what you need. You want to compute the derivative (via product rule)? You want any rel max/min?

OpenStudy (owlcoffee):

study the sign of the derivative.

OpenStudy (anonymous):

in order for obtaining rel max/min?

OpenStudy (owlcoffee):

Yes.

OpenStudy (anonymous):

Sure. The derivative is f '(x) = 3(x-2)(e^(3x-1)) + e^(3x-1) which can be factored as: e^(3x-1) [ 3(x-2) + 1} so f ' (x) = 0 when 3(x-2) + 1 = 0, or x = 5/3.

OpenStudy (owlcoffee):

\[f':f'(x)= e ^{3x+1}+(x-2)(e ^{3x+1})(3)\]

OpenStudy (anonymous):

So, I'll wait till you check that out so that we're on the same page.

OpenStudy (owlcoffee):

So the sign of f'(x) would be positive to the right and negative to the left. meaning that it's a minimum.

OpenStudy (anonymous):

I didnt do that yet..ket me see that for a moment.

OpenStudy (anonymous):

Yes, x = 5/3 will be a rel min.

OpenStudy (owlcoffee):

So f(5/3) = -134.4

OpenStudy (anonymous):

I dont have a cal handy..so I'll take your word for it.

OpenStudy (anonymous):

But (2,0) will be the x-intercept.

OpenStudy (anonymous):

And (0, -2/e) is the y-intercept.

OpenStudy (anonymous):

Which all makes sense, and as x----> + infinity, the graph goes to + infinity, as x = 5/3 is a rel min.

OpenStudy (owlcoffee):

how did you get the intersections?

OpenStudy (anonymous):

To find y - intercept(s), let x = 0. To find x - intercept(s), let y = 0

OpenStudy (anonymous):

work it out.

OpenStudy (owlcoffee):

at y-intercept I got -2e^1

OpenStudy (anonymous):

If x = 0, you get y = (0-2)e^(-1) = -2e^(-1) = -2/e If y = 0, then x = 2 is the only solution (as e^(3x-1) can never equal to 0

OpenStudy (anonymous):

You should have -2e^((-1)

OpenStudy (anonymous):

agree?

OpenStudy (owlcoffee):

yeah, I agree.

OpenStudy (anonymous):

You are plugging in x = 0 into the actual function f(x). Same for y = 0.

OpenStudy (anonymous):

So now you have all the info to sketch the graph.

OpenStudy (owlcoffee):

hmm, I'll need to see the concavity of that, so I'll need to do f''(x)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

You can do that if you want.

OpenStudy (ranga):

BTW, @Owlcoffee, f(5/3) is not -134.4 f(x) = (x - 2)e^(3x - 1) f(5/3) = (5/3 - 2) * e^(3*5/3 - 1) = (-1/3) * e^4 = -18.2